Which pair has equally likely outcomes? List the letters of the two choices below which have equal probabilities of success, separated by a comma. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards).
A. drawing the King of Hearts out of a standard 52 card deck given it’s a face card.
B. rolling a sum of 3 on two fair six sided dice
C. rolling a sum of 5 on two fair six sided dice
D. drawing a queen out of a standard 52 card deck given it’s a face card.
E. rolling a sum of 4 on two fair six sided dice
can some one help me

Answer by ewatrrr(24785)

 
Hi,
standard deck of cards has 12 face cards and four Aces (Aces are not face cards).
A. P(drawing the King of Hearts given it’s a face card.) 1/12
B. P(rolling a sum of 3 on two fair six sided dice) 2/36 = 1/18
C. P(rolling a sum of 5 on two fair six sided dice) = 4/36 = 1/9
D. P(drawing a queen given it’s a face card) = 4/12 = 1/3
E. P(rolling a sum of 4 on two fair six sided dice = 3/36 = 1/12
___1_2_3__4__5__6
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12